in a 19 at an average rate today?
Answer #1

Answer:

Given

rate of increase in cost r=2% per quarter

Number of years =5 years

So number of quarters N=5*4=20

Original cost of development P=$2.1 billion

So Expected cost today A=P*(1+r)^N

A=2.1*(1+2%)^20=$3.12 Billion